OpenStudy (anonymous):

5+3[1+2(2x-3]=6(x+5)

OpenStudy (sandra):

I'm assuming you mean 5+3(1+2(2x-3)) = 6(x+5)

OpenStudy (sandra):

this is another case where you need to distribute

OpenStudy (sandra):

start with the right side, what is 6(x+5) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

6x+30?

OpenStudy (sandra):

yep!

OpenStudy (sandra):

ok now start from the inside out on the left hand side

OpenStudy (anonymous):

do u always start with the right side

OpenStudy (sandra):

so what is 2(2x-3)

OpenStudy (sandra):

well, I think generally start from inside out

OpenStudy (sandra):

as in inner-most thing that needs to be expanded

OpenStudy (anonymous):

4x-6

OpenStudy (sandra):

ok great, so to recap, so far we have: 5 + 3(1 + 4x - 6) = 6x + 30

OpenStudy (anonymous):

gotcha

OpenStudy (sandra):

so to make it easier on yourself, how about combing like terms in the parens, so what is (1+4x - 6) ?

OpenStudy (anonymous):

5x-6 or do u break that down further

OpenStudy (sandra):

well, that's not correct

OpenStudy (sandra):

you can't combine 1 and 4x, since you don't know what 4x

OpenStudy (sandra):

but you can combine 1 and -6

OpenStudy (sandra):

(by adding them)

OpenStudy (anonymous):

-5

OpenStudy (sandra):

if it had been (1x +4x -6) you would have been correct, since the 1 and 4 are both x's

OpenStudy (sandra):

right so the middle is now -5 + 4x, or 4x - 5

OpenStudy (sandra):

so now you have : 5 + 3(4x - 5) = 6x + 30

OpenStudy (sandra):

what is your next step?

OpenStudy (anonymous):

5+12x-15=6x+30

OpenStudy (sandra):

great =)

OpenStudy (sandra):

so can you combine anything on the left side?

OpenStudy (anonymous):

5 and 15 which is -10

OpenStudy (sandra):

yep =), so now what do you have?

OpenStudy (anonymous):

is it 12x-10=6x+30

OpenStudy (sandra):

exactly

OpenStudy (sandra):

so now the basic rule of algebra, and equations/equalities in general - you can perform any operation to both sides of an equation and still have the equatoin be true (satisfied)

OpenStudy (sandra):

the reason that's powerful is because it allows you to eventually get x all by itself

OpenStudy (sandra):

so let's say you wanted to get 12x by itself on the left hand side

OpenStudy (sandra):

what could you add to both sides to make that happen?

OpenStudy (anonymous):

add 10

OpenStudy (sandra):

yep, and what do you get if you add 10 to both sides

OpenStudy (anonymous):

12x=6x+40

OpenStudy (sandra):

exactly. ok , so remember, when you're solving for a variable (in this case x), the idea is to get all of that variable to one side, and then eventually just one instance of that variable by itself (x)

OpenStudy (sandra):

so what could you subtract from both sides, to get all of the x's in one place?

OpenStudy (anonymous):

would u minus 6x

OpenStudy (sandra):

exactly, and then what you get?

OpenStudy (anonymous):

6x=40 then i divided by 6 both sides

OpenStudy (sandra):

yep!

OpenStudy (sandra):

good work

OpenStudy (anonymous):

6.6?

OpenStudy (sandra):

yeah, or you could write it as a fraction

OpenStudy (sandra):

so x = 40/6

OpenStudy (sandra):

and that can be reduced if you divide both top and bottom by 2

OpenStudy (sandra):

so x = 20/3

OpenStudy (sandra):

so now that you think x = 20/3

OpenStudy (anonymous):

thank u so much i think im getting it sum what lol

OpenStudy (sandra):

you stick it back into the original equation (or just use 6.666666 to check in your calculator)

OpenStudy (sandra):

and if the equation is true, then you know you got it right

OpenStudy (sandra):

so for these types of problems, try expanding everything out first

OpenStudy (sandra):

and then just keep doing the same operations on both sides until you conveniently get all the x's (or whatever you're solving for) in one place

OpenStudy (sandra):

and then just get the value for one x by dividing

OpenStudy (sandra):

after a while, you'll start seeing some shortcuts if they're there

OpenStudy (sandra):

for example, if you had 3(x+3) = 12

OpenStudy (sandra):

you wouldn't *have* to expand first

OpenStudy (sandra):

you could just divide both sides by 3

OpenStudy (sandra):

to get x+3 = 4

OpenStudy (sandra):

and then subtract 3 from both sides to get x=1

OpenStudy (sandra):

but seeing that will come with practice

OpenStudy (sandra):

good luck!

OpenStudy (anonymous):

oh ok thanks a bunch !!!!

OpenStudy (sandra):

yep np